find the area bounded by the y² = 2x, y = x - 4.
1
Expert's answer
2011-10-13T10:18:26-0400
Let's transform our system:
x = y2/2, x = y +4;
and change variables:
y = x2/2, y = x +4.
Let’s find the points of interception: x2/2=x+4x2/2-x-4=0 x=4 or x=-2
S= Integral(x+4-x2/2)dx |x=-2 x=4=(x2/2+4x-x3/6) |x=-2 x=4=18